4. The elasticity of demand in the local hardware industry is-1.5, while in the video market it is -4. Which industry has a higher markup over marginal cost (as a percentage of price)? In answering, calculate the markup for each. 5. The local botled water supplier sells water for SO.50 per gallon, while marginal cost is S0.20 per gallon. What is this firms degree of monopoly power, as measured by the Lerner index? 6. What is a network externality? And why might goods that exhibit this externality be more likely to be provided by monopoly suppliers?
Answer #1

Q4) The markup of the local hardware industry can be calculated by using the inverse elasticity rule that is by using the following formula:

(P - MC) / P = 1 / e , where e is the elasticity of demand.

So,    (P - MC) / P = 1 / 1.5

(P - MC) / P = 10 / 15

(P - MC) / P = 0.66

The markup of the video market can be calculated by using the inverse elasticity rule that is by using the following formula:

   (P - MC) / P = 1 / e , where e is the elasticity of demand.

(P - MC) / P = 1 / 4

  (P - MC) / P = 0.25

The local hardware industry has a higher markup over the marginal cost of 0.66 as compared to video market which has a markup of 0.25.

Q5) Price = $0.5

Marginal cost = $ 0.2

The firm's degree of monopoly power can be calculated using the formula of Lerner index that is we can use the following formula:

(P - MC) / P where P is the price charged and MC is the marginal cost of the producer.

= (0.5 - 0.2) / 0.5

= 0.3 / 0.5 = 0.6 is the answer.

Since price is greater than the marginal cost, the firm has some control over the market and the degree of monopoly power is 0.6.